# Hybrid Beamforming Design for Near-Field ISAC with Modular XL-MIMO

### v1

@@ Appendix A Proof of Theorem 1
Therefore, by substituting (65) into (29) and (30), $\mathbf{V}_{c}$ and ${{\mathbf{g}}_{t}}\left({{\mathbf{l}}_{q}}\right),\forall q$ can be expressed as linear transformations of $\mathbf{\tilde{U}}$, namely
${{\mathbf{g}}_{t}}\left({{\mathbf{l}}_{q}}\right)=\mathbf{\tilde{U}}\mathbf{P}^{T}\mathbf{\tilde{{\bm{\nu}}}}_{tq},\ \forall q,$ (66)
@@
${{\mathbf{g}}_{t}}{{\left({{\mathbf{l}}_{q}}\right)}^{H}}{{\mathbf{\tilde{R}}}_{X}}{{\mathbf{g}}_{t}}\left({{\mathbf{l}}_{q}}\right)$ $=\widetilde{\bm{\nu}}_{tq}^{H}\mathbf{P}{{\mathbf{\tilde{U}}}^{H}}{{\mathbf{\tilde{R}}}_{X}}\mathbf{\tilde{U}}{{\mathbf{P}}^{T}}{{\widetilde{\bm{\nu}}}_{tq}}=0,\forall q,$ (68a)
@@
which means that the ${{\mathbf{\tilde{R}}}_{X}}$ component does not contribute to either the communication or sensing performance, but only consumes the transmit power. Thus, we must have $\text{tr}({{{\mathbf{\tilde{R}}}}_{X}})=0$ to satisfy the transmit power constraint. The equality in (70) holds if and only if $\mathbf{\Delta}^{H}\mathbf{P}_{{\tilde{U}}}^{\bot}=\mathbf{0}$, and it can be observed from (63) that this implies ${{\mathbf{\tilde{R}}}_{X}}=\mathbf{0}$.

### v2

@@ Appendix B Proof of Theorem 1
Therefore, by substituting (61) into (24) and (25), $\mathbf{V}_{c}$ and ${{\mathbf{g}}_{tq}},\forall q$ can be expressed as linear transformations of $\mathbf{\tilde{U}}$, namely
${{\mathbf{g}}_{tq}}=\mathbf{\tilde{U}}\mathbf{P}^{T}\mathbf{\tilde{{\bm{\nu}}}}_{tq},\ \forall q,$ (62)
@@
${{\mathbf{g}}_{tq}}^{H}{{\mathbf{\tilde{R}}}_{X}}{{\mathbf{g}}_{tq}})$ $=\widetilde{\bm{\nu}}_{tq}^{H}\mathbf{P}{{\mathbf{\tilde{U}}}^{H}}{{\mathbf{\tilde{R}}}_{X}}\mathbf{\tilde{U}}{{\mathbf{P}}^{T}}{{\widetilde{\bm{\nu}}}_{tq}}=0,\forall q,$ (64a)
@@
which means that the ${{\mathbf{\tilde{R}}}_{X}}$ component does not contribute to either the communication or sensing performance, but only consumes the transmit power[35]. Thus, we must have $\text{tr}({{{\mathbf{\tilde{R}}}}_{X}})=0$ to satisfy the transmit power constraint. The equality in (66) holds if and only if $\mathbf{\Delta}^{H}\mathbf{P}_{{\tilde{U}}}^{\bot}=\mathbf{0}$, and it can be observed from (59) that this implies ${{\mathbf{\tilde{R}}}_{X}}=\mathbf{0}$.
@@
Appendix C Proof of lemma 3
First, we denote $\mathbf{A}\triangleq\mathbf{I}+\frac{1}{\sigma_{c}^{2}}\mathbf{W}_{\text{BB}}^{H}\mathbf{B}\mathbf{W}_{\text{BB}}$. To complete the main proof, we introduce Lemma 4 [36, Theorem 9.B.1].
Lemma 4
Let $\mathbf{R}$ be an $n\times n$ Hermitian matrix with diagonal elements denoted by the vector $\mathbf{d}$ and eigenvalues denoted by the vector $\bm{\lambda}$. Then,
$\mathbf{d}\prec\bm{\lambda},$ (69)
which means $\mathbf{d}$ is majorized by $\bm{\lambda}$[36, Definition 1.A.1].
Since maximization of mutual information is Schur-concave, according to Lemma 4, we have
$\log\det(\bm{\lambda}(\mathbf{A}))\leq\log\det(\text{Diag}(\mathbf{A})),$ (70)
where $\bm{\lambda}(\mathbf{A})$ is the vector of eigenvalues of $\mathbf{A}$ in decreasing order. Therefore, if the optimal $\mathbf{W}_{\text{BB}}^{*}$ does not diagonalize $\mathbf{B}$, we can always find a unitary matrix $\mathbf{M}$ satisfying $\mathbf{M}\mathbf{M}^{H}=\mathbf{I}$, which diagonalizes $\mathbf{B}$. By defining $\tilde{\mathbf{W}}_{\text{BB}}^{*}=\mathbf{W}_{\text{BB}}^{*}\mathbf{M}$, the objective function (35a) is diminished.
It is also noted that
$\operatorname{tr}\left({\tilde{\mathbf{W}}_{\text{BB}}^{*}}(\tilde{\mathbf{W}}_{\text{BB}}^{*})^{H}\right)=\operatorname{tr}\left({\mathbf{W}_{\text{BB}}^{*}}({\mathbf{W}}_{\text{BB}}^{*})^{H}\right)\leq\frac{N_{s}}{M},$ (71a)
$\operatorname{tr}\left((\tilde{\mathbf{W}}_{\text{BB}}^{*})^{H}\bm{\Phi}_{q}\tilde{\mathbf{W}}_{\text{BB}}^{*}\right)=\operatorname{tr}\left(({\mathbf{W}}_{\text{BB}}^{*})^{H}\bm{\Phi}_{q}{\mathbf{W}}_{\text{BB}}^{*}\right),\forall q,$ (71b)
which implies that constraints (35b) and (35c) still hold with $\tilde{\mathbf{W}}_{\text{BB}}^{*}$. This means $\tilde{\mathbf{W}}_{\text{BB}}^{*}$ is a better solution than ${\mathbf{W}}_{\text{BB}}^{*}$, leading to a contradiction. Hence, there exists an optimal solution ${\mathbf{W}}_{\text{BB}}^{*}$ which can diagonalize ${\mathbf{B}}$.
As a result, $\mathbf{A}$ can also be diagonalized by ${\mathbf{W}}_{\text{BB}}^{*}$, yielding ${\mathbf{W}_{\text{BB}}^{*}}^{H}\mathbf{B}\mathbf{W}_{\text{BB}}^{*}=\mathbf{D}$, where $\mathbf{D}$ is a square diagonal matrix. Consequently, we obtain
$\mathbf{B}^{\frac{1}{2}}\mathbf{W}_{\text{BB}}^{*}=\mathbf{\tilde{V}}\tilde{\bm{\Sigma}},$ (72)
where $\mathbf{\tilde{V}}$ is an arbitrary unitary matrix, $\bm{\Sigma}=\left[\begin{array}[]{c}\mathbf{D}^{1/2}\\ \mathbf{0}\end{array}\right]$ and $\bm{\Sigma}^{H}\bm{\Sigma}=\mathbf{D}$ Taking SVD of $\mathbf{B}$, we have $\mathbf{B}={{\mathbf{U}}_{\text{B}}}\mathbf{\Sigma}_{\text{B}}\mathbf{U}_{\text{B}}^{H}$. Based on (72), it follows that
$\mathbf{W}_{\text{BB}}^{*}=\mathbf{B}^{-\frac{1}{2}}\mathbf{\tilde{V}}\tilde{\bm{\Sigma}}={{\mathbf{U}}_{\text{B}}}\mathbf{\Sigma}_{\text{B}}^{-\frac{1}{2}}\mathbf{U}_{\text{B}}^{H}\mathbf{\tilde{V}}\tilde{\bm{\Sigma}},$ (73)
which completes the proof.